# Retrieval Pivot Attacks in Hybrid RAG: Measuring and Mitigating Amplified Leakage from Vector Seeds to Graph Expansion

### v1

@@ Abstract
Hybrid Retrieval-Augmented Generation (RAG) pipelines increasingly combine vector similarity search with knowledge graph expansion to support multi-hop reasoning. We show this composition introduces a distinct security failure mode: a semantically retrieved “seed” chunk can pivot via entity linking into sensitive graph neighborhoods, causing data leakage that does not exist in vector-only retrieval. We formalize this risk as Retrieval Pivot Risk (RPR) and define companion metrics Leakage@k, Amplification Factor (AF), and Pivot Depth (PD) to quantify leakage probability, magnitude, amplification over vector baselines, and structural distance to the first unauthorized node. We further present a taxonomy of four Retrieval Pivot Attacks that exploit the vector-to-graph boundary with injection budgets as small as 10–20 chunks, and we demonstrate that adversarial injection is not required: naturally shared entities (e.g., vendors, infrastructure, compliance standards) create cross-tenant pivot paths organically. In a synthetic multi-tenant enterprise corpus (1,000 documents; 2,785 graph nodes; 15,514 edges) evaluated with 500 queries, the undefended hybrid pipeline exhibits $\text{RPR}\approx 0.95$ and $\text{AF}(\epsilon)\approx 160\text{--}194\times$ relative to vector-only retrieval, with leakage occurring at $\text{PD}=2$ hops as a structural consequence of the bipartite chunk–entity topology. We validate these findings on the Enron email corpus (50,000 emails; 376,000 graph nodes; 2.3M edges), where undefended hybrid retrieval produces $\text{RPR}=0.695$ on benign queries—lower than the synthetic corpus due to sparser entity connectivity, but still affecting 70% of queries with a mean of 7.1 leaked items each. We propose five layered defenses and find that a single placement fix—per-hop authorization at the graph expansion boundary—eliminates all measured leakage ($\text{RPR}\to 0.0$) across both corpora, all queries, and all attack variants with negligible latency overhead, indicating the vulnerability is primarily a boundary enforcement problem rather than a defense complexity problem.
I Introduction
@@
Enterprise adoption of Retrieval-Augmented Generation (RAG) has accelerated rapidly: 30–60% of enterprise AI use cases now rely on RAG architectures [9], and the vector database market reached \$1.73 billion in 2024 [14]. To improve multi-hop reasoning over complex organizational knowledge, practitioners increasingly deploy hybrid RAG pipelines that combine vector similarity search with knowledge graph expansion [16, 15]. These systems retrieve an initial set of text chunks via embedding similarity, then pivot through entity mentions into a knowledge graph to gather structurally related context before passing the assembled results to a large language model (LLM). We call the transition between vector retrieval and graph expansion the pivot boundary: the architectural point where a retrieved seed becomes a graph traversal capability.
@@
Prior work falls into two separate buckets that do not address this boundary. (1) Vector-only attacks: PoisonedRAG achieves 90% attack success rate by injecting 5 malicious texts into vector stores [28]; CorruptRAG [24] and CtrlRAG [18] extend these results. (2) Graph-only attacks: GRAGPoison demonstrates 98% success via relation-centric poisoning within GraphRAG [8]; TKPA [20] and RAG-Safety [25] target graph-side integrity. (3) The hybrid boundary—how vector outputs become graph seeds—has not been systematically measured or formalized. The OWASP LLM Top 10 (2025) introduced LLM08 (Vector and Embedding Weaknesses) but does not address graph components [13]. MITRE ATLAS treats RAG as a monolithic system [11]. The SoK on RAG privacy explicitly identifies hybrid RAG security as an open problem [2].
@@
    Composition vulnerability + metrics. We formalize Retrieval Pivot Risk (RPR) and companion metrics (AF, PD, Leakage@k, severity-weighted leakage) that quantify a compound attack surface emerging from composing two individually secure retrieval modalities. In our bipartite chunk–entity graph, all leakage occurs at PD == 2 hops—a structural signature of the pivot boundary (§V).
@@
    Cross-dataset validation + organic leakage. We demonstrate the vulnerability on both a synthetic enterprise corpus (1,000 documents, 4 tenants) and the Enron email corpus (50,000 emails, 5 departments), showing RPR == 0.95 and RPR == 0.70 respectively without adversarial injection. We present four Retrieval Pivot Attacks (A1–A4) exploiting the pivot boundary with injection budgets of 10–20 chunks (§IV, §VII).
@@
    Defense placement analysis. We evaluate five layered defenses (D1–D5) and show that authorization must be re-checked at the graph expansion boundary. D1 (per-hop authorization) alone eliminates all measured leakage ($\text{RPR}\to 0.0$) across both corpora, all queries, all attack variants, and metadata mislabel rates up to 5%. D2–D5 serve as utility optimizers and defense-in-depth layers (§VIII, §VII).
@@
II Background
II-A Vector Retrieval in RAG
Standard RAG systems encode documents as dense vector embeddings using models such as all-MiniLM-L6-v2 [26] and retrieve the top-$k$ chunks by cosine similarity to the query embedding. In multi-tenant deployments, vector stores apply metadata prefilters (e.g., tenant ID) before similarity search, ensuring that retrieval respects organizational boundaries. This prefiltering makes vector-only RAG robust against cross-tenant leakage: our experiments confirm $\text{RPR}=0.0$ for vector-only pipelines across all evaluation queries on both synthetic and Enron corpora (§VII).
II-B Knowledge Graph RAG
GraphRAG [15] and related approaches construct knowledge graphs from document corpora via named entity recognition (NER) and relation extraction, then leverage graph structure for multi-hop reasoning. Systems like AgCyRAG [7] use LLM-driven graph traversal for complex queries. Graph-based retrieval provides superior comprehensiveness on multi-hop queries—86% versus 57% for vector RAG [16]—but introduces graph-specific attack surfaces: GRAGPoison [8] achieves 98% attack success through relation-centric poisoning, and TKPA [20] drops QA accuracy from 95% to 50% by modifying just 0.06% of corpus text.
II-C Hybrid RAG and the Pivot Boundary
@@
III Threat Model
III-A System Model
@@ III-B Attacker Capabilities
Injection attacker. The attacker holds a legitimate account in one tenant (acme_engineering) and can inject documents into the shared corpus through standard channels—wiki edits, ticket updates, shared document repositories. Injected documents undergo normal ingestion: chunking, NER-based entity extraction, embedding, and indexing into both vector and graph stores. The attacker’s injection budget is modest: 10–20 chunks, consistent with the budgets shown effective in prior work [28, 24].
No-injection attacker. The attacker holds a legitimate account and crafts queries that mention bridge entities (shared infrastructure names, vendor references, cross-team personnel) to maximize pivot probability through naturally occurring cross-tenant entity connections. This attacker requires zero document injection—the organic structure of the multi-tenant knowledge graph provides the pivot paths. Our evaluation shows that this attacker model achieves $\text{RPR}=0.95$ through benign queries alone (§VII), demonstrating that the vulnerability is structural rather than injection-dependent.
III-C Attacker Goals
@@
III-D Adaptive Attacker Considerations
An adaptive attacker aware of D1 (per-hop authorization) might attempt to bypass it through: (a) metadata spoofing—forging tenant or sensitivity labels during document injection, or (b) same-tenant escalation—crafting documents within the attacker’s own tenant that link to higher-sensitivity content within the same tenant boundary. D1 is robust against (a) if and only if the ingestion pipeline enforces metadata integrity (the metadata is assigned by the system, not the uploader). We evaluate D1’s robustness under metadata corruption in §VII-K and discuss mitigation strategies in §VIII-B. Same-tenant escalation (b) targets sensitivity boundaries rather than tenant boundaries and requires the attacker to already have write access to the target tenant’s corpus.
III-E Metadata Integrity Assumption
D1 (per-hop authorization) relies on node metadata—tenant labels and sensitivity tiers—being trustworthy. We assume metadata is system-assigned during document ingestion (the uploader cannot choose tenant or sensitivity labels). If an attacker can forge metadata, D1 can be bypassed. We evaluate robustness under random metadata corruption in §VII-K and discuss additional integrity mechanisms in §VIII-B.
III-F Scope
We measure unauthorized items in the retrieval context window, not exfiltrated tokens or generated text. We focus exclusively on the retrieval layer. We do not evaluate LLM generation quality, jailbreaking, or prompt injection. The attacker does not have direct database access, cannot modify model weights, and cannot alter the pipeline configuration. We omit a graph-only baseline (P2) because graph-only RAG without vector seeding is a fundamentally different retrieval paradigm—our question is amplification created by composition, not graph-only security, which is addressed by prior work [8, 20].
IV Retrieval Pivot Attacks
We define four non-adaptive attacks (A1–A4) that exploit the pivot boundary with increasing structural sophistication, and three adaptive attacks (A5–A7) that target the defense mechanisms themselves. Each attack operates within the threat model of §III: the attacker injects a small number of crafted documents that undergo standard ingestion.
IV-A A1: Seed Steering
@@
IV-B A2: Entity Anchor Injection
@@ IV-C A3: Neighborhood Flooding
IV-D A4: Bridge Node Attack
@@ V Metrics
Item serialization. Each item in $S_{k}(q)$ is a node serialized for the LLM context: chunk nodes are serialized as their source text; entity nodes are serialized as {name, type, top-3 relations}. Context size counts the total number of serialized nodes (chunks + entities) presented to the LLM. The default pipeline configuration (depth $d=2$, branching $\leq 10$, total $\leq 100$ nodes, top-k=10k{=}10 vector results) produces $\sim$}110 items per query in the undefended hybrid pipeline (P3). Table I summarizes the notation used throughout.
TABLE I: Notation summary.
@@
V-A Retrieval Pivot Risk (RPR)
@@ V-B Leakage@k
V-C Severity-Weighted Leakage
@@ V-E Pivot Depth (PD)
VI Experimental Setup
VI-A Synthetic Enterprise Corpus
@@
VI-B Knowledge Graph Construction
@@
VI-C Enron Email Corpus
@@
VI-D Pipeline Variants
We evaluate 7 pipeline configurations (Table II):
TABLE II: Pipeline variants and their defense configurations.
@@ VI-D Pipeline Variants
We omit a graph-only baseline (P2) because graph-only retrieval without vector seeding is a fundamentally different paradigm that does not involve the pivot boundary under study. Its security properties are addressed by prior work on GraphRAG attacks [8, 20].
VI-E Evaluation Protocol
@@ VI-F Attack Evaluation
VII Results
VII-A Hybrid RAG Amplifies Leakage
Table III presents security metrics across all pipeline variants. The central finding is stark: the undefended hybrid pipeline (P3) leaks massively while the vector-only baseline (P1) leaks nothing.
@@ VII-A Hybrid RAG Amplifies Leakage
TABLE III: Defense ablation: security metrics across pipeline variants (500 queries: 350 benign + 150 adversarial). RPR = Retrieval Pivot Risk with 95% bootstrap CI, Leak = mean Leakage@k, SWL = severity-weighted leakage, PD = Pivot Depth (hops), Ctx = mean context size. “–” indicates no leakage occurred (PD undefined).
@@ VII-B The PD = 2 Structural Signature
This 2-hop pattern is inherent to any hybrid RAG system that constructs a bipartite graph between chunks and entities with expansion depth $d\geq 2$: the entity-linking step creates a structural bridge between vector-retrieved content and graph-stored content. The PD=2 finding is structural in our bipartite construction—knowledge graphs with richer entity-to-entity relationships (e.g., ontological hierarchies, multi-hop inference chains) may exhibit leakage at deeper pivot depths. We discuss how PD varies with traversal parameters in §VII-I.
VII-C Organic Leakage: No Injection Required
@@
Bridge category analysis. To understand which types of shared entities drive leakage, we analyze the hop-1 pivot nodes in each leaking query’s traversal path and classify them against the 5 bridge entity categories (Table IV).
TABLE IV: Organic leakage by bridge entity category under P3 (no injection). Queries = leaking queries with that bridge type at hop 1. Leak = attributed leaked items.
@@
VII-D Connectivity Sensitivity
To measure how shared-entity density affects leakage, we regenerate the corpus with bridge entity counts $\in\{0,5,10,15,25,40\}$, rebuild all indexes for each count, and run 100 adversarial queries through P3 (Table V).
TABLE V: Connectivity sweep: effect of bridge entity count on P3 RPR and mean Leakage@k. RPR is remarkably stable; mean leakage increases monotonically with connectivity.

### v2

@@ Abstract
Hybrid Retrieval-Augmented Generation (RAG) pipelines increasingly combine vector similarity search with knowledge graph expansion to support multi-hop reasoning. We show this composition introduces a distinct security failure mode: a semantically retrieved “seed” chunk can pivot via entity linking into sensitive graph neighborhoods, causing data leakage that does not exist in vector-only retrieval. We formalize this risk as Retrieval Pivot Risk (RPR) and define companion metrics Leakage@k, Amplification Factor (AF), and Pivot Depth (PD) to quantify leakage probability, magnitude, amplification over vector baselines, and structural distance to the first unauthorized node. We further present a taxonomy of four Retrieval Pivot Attacks that exploit the vector-to-graph boundary with injection budgets as small as 10–20 chunks, and we demonstrate that adversarial injection is not required: naturally shared entities (e.g., vendors, infrastructure, compliance standards) create cross-tenant pivot paths organically. In a synthetic multi-tenant enterprise corpus (1,000 documents; 2,785 graph nodes; 15,514 edges) evaluated with 500 queries, the undefended hybrid pipeline exhibits $\text{RPR}\approx 0.95$ and $\text{AF}(\epsilon)\approx 160\text{--}194\times$ relative to vector-only retrieval, with leakage occurring at $\text{PD}=2$ hops as a structural consequence of the bipartite chunk–entity topology. We validate these findings on two real-world corpora: the Enron email corpus (50,000 emails; $\text{RPR}=0.695$) and SEC EDGAR 10-K filings (887 sections across 20 companies; $\text{RPR}=0.085$). RPR scales with entity connectivity density, but the structural invariant ($\text{PD}=2$) persists across all three corpora. We propose five layered defenses and find that a single placement fix—per-hop authorization at the graph expansion boundary—eliminates all measured leakage ($\text{RPR}\to 0.0$) across all three corpora, all queries, and all attack variants with negligible latency overhead, indicating the vulnerability is primarily a boundary enforcement problem rather than a defense complexity problem.
1 Introduction
@@
Enterprise adoption of Retrieval-Augmented Generation (RAG) has accelerated rapidly: 30–60% of enterprise AI use cases now rely on RAG architectures [saferag2025], and the vector database market reached \$1.73 billion in 2024 [pineconerac2024]. To improve multi-hop reasoning over complex organizational knowledge, practitioners increasingly deploy hybrid RAG pipelines that combine vector similarity search with knowledge graph expansion [hybridrag2024, microsoftgraphrag2024]. These systems retrieve an initial set of text chunks via embedding similarity, then pivot through entity mentions into a knowledge graph to gather structurally related context before passing the assembled results to a large language model (LLM). We call the transition between vector retrieval and graph expansion the pivot boundary: the architectural point where a retrieved seed becomes a graph traversal capability.
@@
Prior work falls into two separate buckets that do not address this boundary. (1) Vector-only attacks: PoisonedRAG achieves 90% attack success rate by injecting 5 malicious texts into vector stores [zou2025poisonedrag]; CorruptRAG [corruptrag2025] and CtrlRAG [ctrlrag2025] extend these results. (2) Graph-only attacks: GRAGPoison demonstrates 98% success via relation-centric poisoning within GraphRAG [gragpoison2026]; TKPA [fewwords2025] and RAG-Safety [ragsafety2025] target graph-side integrity. (3) The hybrid boundary—how vector outputs become graph seeds—has not been systematically measured or formalized. The OWASP LLM Top 10 (2025) introduced LLM08 (Vector and Embedding Weaknesses) but does not address graph components [owasptop10llm2025]. MITRE ATLAS treats RAG as a monolithic system [mitreatlas2025]. The SoK on RAG privacy explicitly identifies hybrid RAG security as an open problem [sokprivacy2026].
@@
    Composition vulnerability + metrics. We formalize Retrieval Pivot Risk (RPR) and companion metrics (AF, PD, Leakage@k, severity-weighted leakage) that quantify a compound attack surface emerging from composing two individually secure retrieval modalities. In our bipartite chunk–entity graph, all leakage occurs at PD == 2 hops—a structural signature of the pivot boundary (§5).
@@
    Cross-dataset validation + organic leakage. We demonstrate the vulnerability on three corpora: a synthetic enterprise corpus (1,000 documents, 4 tenants; RPR == 0.95), the Enron email corpus (50,000 emails, 5 departments; RPR == 0.70), and SEC EDGAR 10-K filings (887 sections, 4 sectors; RPR == 0.09)—all without adversarial injection. We present four Retrieval Pivot Attacks (A1–A4) exploiting the pivot boundary with injection budgets of 10–20 chunks (§4, §7).
@@
    Defense placement analysis. We evaluate five layered defenses (D1–D5) and show that authorization must be re-checked at the graph expansion boundary. D1 (per-hop authorization) alone eliminates all measured leakage ($\text{RPR}\to 0.0$) across both corpora, all queries, all attack variants, and metadata mislabel rates up to 5%. D2–D5 serve as utility optimizers and defense-in-depth layers (§8, §7).
@@
2 Background
2.1 Vector Retrieval in RAG
Standard RAG systems encode documents as dense vector embeddings using models such as all-MiniLM-L6-v2 [corpuspoisoning2023] and retrieve the top-$k$ chunks by cosine similarity to the query embedding. In multi-tenant deployments, vector stores apply metadata prefilters (e.g., tenant ID) before similarity search, ensuring that retrieval respects organizational boundaries. This prefiltering makes vector-only RAG robust against cross-tenant leakage: our experiments confirm $\text{RPR}=0.0$ for vector-only pipelines across all evaluation queries on all three corpora (§7).
2.2 Knowledge Graph RAG
GraphRAG [microsoftgraphrag2024] and related approaches construct knowledge graphs from document corpora via named entity recognition (NER) and relation extraction, then leverage graph structure for multi-hop reasoning. Systems like AgCyRAG [agcyrag2025] use LLM-driven graph traversal for complex queries. Graph-based retrieval provides superior comprehensiveness on multi-hop queries—86% versus 57% for vector RAG [hybridrag2024]—but introduces graph-specific attack surfaces: GRAGPoison [gragpoison2026] achieves 98% attack success through relation-centric poisoning, and TKPA [fewwords2025] drops QA accuracy from 95% to 50% by modifying just 0.06% of corpus text.
2.3 Hybrid RAG and the Pivot Boundary
@@
3 Threat Model
3.1 System Model
@@ 3.2 Attacker Capabilities
Injection attacker. The attacker holds a legitimate account in one tenant (acme_engineering) and can inject documents into the shared corpus through standard channels—wiki edits, ticket updates, shared document repositories. Injected documents undergo normal ingestion: chunking, NER-based entity extraction, embedding, and indexing into both vector and graph stores. The attacker’s injection budget is modest: 10–20 chunks, consistent with the budgets shown effective in prior work [zou2025poisonedrag, corruptrag2025].
No-injection attacker. The attacker holds a legitimate account and crafts queries that mention bridge entities (shared infrastructure names, vendor references, cross-team personnel) to maximize pivot probability through naturally occurring cross-tenant entity connections. This attacker requires zero document injection—the organic structure of the multi-tenant knowledge graph provides the pivot paths. Our evaluation shows that this attacker model achieves $\text{RPR}=0.95$ through benign queries alone (§7), demonstrating that the vulnerability is structural rather than injection-dependent.
3.3 Attacker Goals
@@
3.4 Adaptive Attacker Considerations
An adaptive attacker aware of D1 (per-hop authorization) might attempt to bypass it through: (a) metadata spoofing—forging tenant or sensitivity labels during document injection, or (b) same-tenant escalation—crafting documents within the attacker’s own tenant that link to higher-sensitivity content within the same tenant boundary. D1 is robust against (a) if and only if the ingestion pipeline enforces metadata integrity (the metadata is assigned by the system, not the uploader). We evaluate D1’s robustness under metadata corruption in §7.11 and discuss mitigation strategies in §8.2. Same-tenant escalation (b) targets sensitivity boundaries rather than tenant boundaries and requires the attacker to already have write access to the target tenant’s corpus.
3.5 Metadata Integrity Assumption
D1 (per-hop authorization) relies on node metadata—tenant labels and sensitivity tiers—being trustworthy. We assume metadata is system-assigned during document ingestion (the uploader cannot choose tenant or sensitivity labels). If an attacker can forge metadata, D1 can be bypassed. We evaluate robustness under random metadata corruption in §7.11 and discuss additional integrity mechanisms in §8.2.
3.6 Scope
We measure unauthorized items in the retrieval context window, not exfiltrated tokens or generated text. We focus exclusively on the retrieval layer. We do not evaluate LLM generation quality, jailbreaking, or prompt injection. The attacker does not have direct database access, cannot modify model weights, and cannot alter the pipeline configuration. We omit a graph-only baseline (P2) because graph-only RAG without vector seeding is a fundamentally different retrieval paradigm—our question is amplification created by composition, not graph-only security, which is addressed by prior work [gragpoison2026, fewwords2025].
4 Retrieval Pivot Attacks
We define four non-adaptive attacks (A1–A4) that exploit the pivot boundary with increasing structural sophistication, and three adaptive attacks (A5–A7) that target the defense mechanisms themselves. Each attack operates within the threat model of §3: the attacker injects a small number of crafted documents that undergo standard ingestion.
4.1 A1: Seed Steering
@@
4.2 A2: Entity Anchor Injection
@@ 4.3 A3: Neighborhood Flooding
4.4 A4: Bridge Node Attack
@@ 5 Metrics
Item serialization. Each item in $S_{k}(q)$ is a node serialized for the LLM context: chunk nodes are serialized as their source text; entity nodes are serialized as {name, type, top-3 relations}. Context size counts the total number of serialized nodes (chunks + entities) presented to the LLM. The default pipeline configuration (depth $d=2$, branching $\leq 10$, total $\leq 100$ nodes, top-k=10k{=}10 vector results) produces $\sim$}110 items per query in the undefended hybrid pipeline (P3). Table 1 summarizes the notation used throughout.
Table 1: Notation summary.
@@
5.1 Retrieval Pivot Risk (RPR)
@@ 5.2 Leakage@k
5.3 Severity-Weighted Leakage
@@ 5.5 Pivot Depth (PD)
6 Experimental Setup
6.1 Synthetic Enterprise Corpus
@@
6.2 Knowledge Graph Construction
@@
6.3 Enron Email Corpus
@@
6.4 EDGAR 10-K Corpus
As a third corpus, we evaluate on SEC EDGAR 10-K annual reports—public filings from 20 companies across four industry sectors (tech, finance, healthcare, energy). Each sector serves as a tenant, and filing sections receive sensitivity labels based on a Material Non-Public Information (MNPI) framework: Items 1–4 (public), Items 7–8 (internal), Items 1A/7A/9A/11–12 (confidential), Items 5/13 (restricted).
The EDGAR graph contains 19,527 nodes (887 documents, 3,692 chunks, 14,948 entities) and 427,415 edges. Cross-sector bridge entities arise naturally from shared auditors (Big 4 firms), institutional investors (BlackRock, Vanguard), and board members serving on companies in multiple sectors. NER extraction on formal financial language produces dense entity graphs ($\sim$4 entity mentions per chunk) but the standardized vocabulary creates fewer distinctive cross-sector entity connections than either the synthetic or Enron corpus.
Key differences from other corpora. EDGAR filings use formal, standardized language with fewer organic entity connections across sector-tenants than Enron’s informal email traffic. This produces the lowest RPR of the three corpora ($0.085$ vs. $0.695$ Enron, $0.954$ synthetic) while preserving the $\text{PD}=2$ structural invariant.
6.5 Pipeline Variants
We evaluate 7 pipeline configurations (Table 2):
Table 2: Pipeline variants and their defense configurations.
@@ 6.5 Pipeline Variants
We omit a graph-only baseline (P2) because graph-only retrieval without vector seeding is a fundamentally different paradigm that does not involve the pivot boundary under study. Its security properties are addressed by prior work on GraphRAG attacks [gragpoison2026, fewwords2025].
6.6 Evaluation Protocol
@@ 6.7 Attack Evaluation
7 Results
7.1 Hybrid RAG Amplifies Leakage
Table 3 presents security metrics across all pipeline variants. The central finding is stark: the undefended hybrid pipeline (P3) leaks massively while the vector-only baseline (P1) leaks nothing.
@@ 7.1 Hybrid RAG Amplifies Leakage
Table 3: Defense ablation: security metrics across pipeline variants (500 queries: 350 benign + 150 adversarial). RPR = Retrieval Pivot Risk with 95% bootstrap CI, Leak = mean Leakage@k, SWL = severity-weighted leakage, PD = Pivot Depth (hops), Ctx = mean context size. “–” indicates no leakage occurred (PD undefined).
@@ 7.2 The PD = 2 Structural Signature
This 2-hop pattern is inherent to any hybrid RAG system that constructs a bipartite graph between chunks and entities with expansion depth $d\geq 2$: the entity-linking step creates a structural bridge between vector-retrieved content and graph-stored content. The PD=2 finding is structural in our bipartite construction—knowledge graphs with richer entity-to-entity relationships (e.g., ontological hierarchies, multi-hop inference chains) may exhibit leakage at deeper pivot depths. We discuss how PD varies with traversal parameters in §7.9.
7.3 Organic Leakage: No Injection Required
@@
Bridge category analysis. To understand which types of shared entities drive leakage, we analyze the hop-1 pivot nodes in each leaking query’s traversal path and classify them against the 5 bridge entity categories (Table 4).
Table 4: Organic leakage by bridge entity category under P3 (no injection). Queries = leaking queries with that bridge type at hop 1. Leak = attributed leaked items.
@@
7.4 Connectivity Sensitivity
To measure how shared-entity density affects leakage, we regenerate the corpus with bridge entity counts $\in\{0,5,10,15,25,40\}$, rebuild all indexes for each count, and run 100 adversarial queries through P3 (Table 5).
Table 5: Connectivity sweep: effect of bridge entity count on P3 RPR and mean Leakage@k. RPR is remarkably stable; mean leakage increases monotonically with connectivity.